For what values of ๐‘ does the curve ๐‘“(๐‘ฅ) = 2๐‘ฅ^3 + ๐‘๐‘ฅ^2 + 2๐‘ฅ have maximum and minimum points?









Show that the minimum and maximum points of every curve in the family of polynomials







๐‘“(๐‘ฅ) = 2๐‘ฅ^3 + ๐‘๐‘ฅ^2 + 2๐‘ฅ lie on the curve ๐‘ฆ = ๐‘ฅ โˆ’ ๐‘ฅ^3.














Suppose ๐‘“ is differentiable on โ„ and has two roots. Show that ๐‘“โ€ฒ has at least one root.

1
Expert's answer
2022-05-09T18:20:29-0400

1.


"f'(x)=6x^2+2cx+2"

Find the critical number(s)


"f'(x)=0=>6x^2+2cx+2=0"

"3x^2+cx+1=0""x=\dfrac{-c\pm\sqrt{c^2-12}}{6}"

We consider "x\in \R"

"c^2-12\ge0=>c\le-\sqrt{12}\ or\ c\ge\sqrt{12}"

The curve "\ud835\udc53(\ud835\udc65) = 2\ud835\udc65^3 + \ud835\udc50\ud835\udc65^2 + 2\ud835\udc65" have maximum and minimum for "c\in(-\infin, -\sqrt{12})\cup (\sqrt{12}, \infin)."


2.



"f'(x)=6x^2+2cx+2"

Find the critical number(s)


"f'(x)=0=>6x^2+2cx+2=0"

"3x^2+cx+1=0""x=\dfrac{-c\pm\sqrt{c^2-12}}{6}"

We consider "x\in \R"

"c^2-12\ge0=>c\le-\sqrt{12}\ or\ c\ge\sqrt{12}"

In this case


"x\not=0, cx^2=-3x^3-x"

Substitute


"y = 2x^3-3x^3-x+2x"

"y = -x^3+x"

Then the minimum and maximum points of every curve in the family of polynomialsย "f(x)=2x^3+cx^2+2x" lie on the curve "y=x-x^3."

"x\not=0, c\le-\sqrt{12}" or "c\ge\sqrt{12}."


3.

We say that "f" has roots at "a" and "b, (a<b)," and since "f" is continuous and differentiable everywhere as polynomial, it is continuous on "[a, b]" and differentiable on "(a, b)." So by the Rolleโ€™s theorem, there is a number "k" in "(a, b)," such that "f'(k)=0." Therefore, "k" is a root of "f'(x)=0."